# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

### v1

@@ Appendix G Full Results over Problems
QwQ-32B $0.94_{0.90}^{1.00}$ $0.87_{0.83}^{0.93}$ $0.80_{0.70}^{0.93}$ $0.62_{0.57}^{0.67}$ $0.33_{0.23}^{0.40}$ $0.16_{0.10}^{0.20}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.69_{0.67}^{0.73}$ $0.36_{0.27}^{0.40}$ $0.24_{0.13}^{0.40}$ $0.09_{0.03}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.70_{0.67}^{0.73}$ $0.26_{0.13}^{0.40}$ $0.09_{0.07}^{0.10}$ $0.08_{0.03}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.73_{0.73}^{0.73}$ $0.39_{0.33}^{0.43}$ $0.22_{0.17}^{0.27}$ $0.12_{0.07}^{0.20}$ $0.09_{0.00}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.06_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.99_{0.97}^{1.00}$ $0.80_{0.70}^{0.90}$ $0.74_{0.67}^{0.83}$ $0.64_{0.53}^{0.77}$ $0.32_{0.17}^{0.50}$ $0.23_{0.20}^{0.27}$ $0.27_{0.20}^{0.33}$ $0.16_{0.10}^{0.27}$ $0.10_{0.10}^{0.10}$ $0.02_{0.00}^{0.07}$
DeepSeek-V3 $0.83_{0.77}^{0.90}$ $0.44_{0.30}^{0.53}$ $0.14_{0.10}^{0.20}$ $0.16_{0.13}^{0.17}$ $0.09_{0.00}^{0.17}$ $0.06_{0.03}^{0.07}$ $0.06_{0.00}^{0.10}$ $0.06_{0.00}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$
DeepSeek-V3-2503 $0.99_{0.97}^{1.00}$ $0.82_{0.77}^{0.90}$ $0.51_{0.50}^{0.53}$ $0.38_{0.27}^{0.53}$ $0.16_{0.07}^{0.27}$ $0.14_{0.13}^{0.17}$ $0.09_{0.07}^{0.10}$ $0.10_{0.07}^{0.17}$ $0.06_{0.03}^{0.07}$ $0.03_{0.00}^{0.07}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.91_{0.83}^{0.97}$ $0.76_{0.63}^{0.93}$ $0.64_{0.57}^{0.73}$ $0.49_{0.43}^{0.57}$ $0.36_{0.27}^{0.40}$ $0.17_{0.13}^{0.23}$ $0.04_{0.00}^{0.10}$
o1-mini $0.72_{0.57}^{0.80}$ $0.71_{0.67}^{0.77}$ $0.54_{0.50}^{0.60}$ $0.40_{0.33}^{0.47}$ $0.19_{0.17}^{0.23}$ $0.23_{0.13}^{0.30}$ $0.12_{0.10}^{0.17}$ $0.08_{0.03}^{0.10}$ $0.04_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$
o3-mini $0.82_{0.80}^{0.83}$ $0.84_{0.77}^{0.90}$ $0.71_{0.63}^{0.77}$ $0.71_{0.60}^{0.83}$ $0.63_{0.57}^{0.73}$ $0.59_{0.50}^{0.67}$ $0.44_{0.37}^{0.50}$ $0.32_{0.27}^{0.43}$ $0.20_{0.10}^{0.33}$ $0.22_{0.20}^{0.23}$
@@
QwQ-32B $0.88_{0.80}^{0.93}$ $0.83_{0.77}^{0.87}$ $0.46_{0.40}^{0.57}$ $0.08_{0.00}^{0.20}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.26_{0.17}^{0.33}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.30_{0.23}^{0.33}$ $0.03_{0.03}^{0.03}$ $0.04_{0.00}^{0.10}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.83_{0.73}^{0.90}$ $0.44_{0.40}^{0.50}$ $0.34_{0.33}^{0.37}$ $0.18_{0.13}^{0.20}$ $0.04_{0.00}^{0.10}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.98_{0.97}^{1.00}$ $0.89_{0.83}^{0.93}$ $0.58_{0.43}^{0.70}$ $0.39_{0.37}^{0.40}$ $0.07_{0.00}^{0.17}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.66_{0.60}^{0.73}$ $0.46_{0.40}^{0.50}$ $0.44_{0.37}^{0.50}$ $0.37_{0.33}^{0.40}$ $0.06_{0.00}^{0.13}$ $0.04_{0.03}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.87_{0.80}^{0.93}$ $0.74_{0.67}^{0.83}$ $0.62_{0.57}^{0.67}$ $0.18_{0.10}^{0.27}$ $0.18_{0.13}^{0.23}$ $0.09_{0.03}^{0.13}$ $0.02_{0.00}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.80_{0.73}^{0.90}$ $0.64_{0.57}^{0.77}$ $0.49_{0.47}^{0.53}$ $0.29_{0.20}^{0.43}$ $0.06_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$
o1-mini $0.67_{0.43}^{0.80}$ $0.58_{0.50}^{0.63}$ $0.52_{0.47}^{0.57}$ $0.33_{0.27}^{0.40}$ $0.31_{0.20}^{0.50}$ $0.19_{0.13}^{0.23}$ $0.07_{0.03}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.02_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$
@@
QwQ-32B $0.96_{0.90}^{1.00}$ $0.91_{0.87}^{0.93}$ $0.78_{0.73}^{0.87}$ $0.56_{0.50}^{0.67}$ $0.34_{0.27}^{0.43}$ $0.10_{0.07}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.49_{0.43}^{0.57}$ $0.51_{0.43}^{0.57}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.40_{0.30}^{0.50}$ $0.17_{0.13}^{0.20}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.60_{0.57}^{0.63}$ $0.39_{0.30}^{0.53}$ $0.03_{0.00}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.76_{0.67}^{0.87}$ $0.70_{0.67}^{0.73}$ $0.22_{0.07}^{0.33}$ $0.17_{0.13}^{0.20}$ $0.09_{0.07}^{0.10}$ $0.04_{0.03}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.67_{0.63}^{0.70}$ $0.60_{0.53}^{0.63}$ $0.13_{0.07}^{0.20}$ $0.12_{0.10}^{0.17}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.07}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $0.80_{0.73}^{0.87}$ $0.90_{0.87}^{0.93}$ $0.48_{0.40}^{0.63}$ $0.64_{0.57}^{0.73}$ $0.32_{0.30}^{0.37}$ $0.16_{0.07}^{0.20}$ $0.16_{0.07}^{0.23}$ $0.09_{0.00}^{0.20}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $0.99_{0.97}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.97_{0.97}^{0.97}$ $0.88_{0.80}^{0.93}$ $0.72_{0.67}^{0.77}$ $0.72_{0.67}^{0.80}$ $0.51_{0.40}^{0.67}$ $0.22_{0.17}^{0.27}$ $0.04_{0.03}^{0.07}$
o1-mini $0.61_{0.50}^{0.70}$ $0.76_{0.70}^{0.87}$ $0.57_{0.43}^{0.70}$ $0.62_{0.60}^{0.67}$ $0.37_{0.33}^{0.43}$ $0.27_{0.23}^{0.30}$ $0.34_{0.27}^{0.40}$ $0.17_{0.07}^{0.23}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.07}$
@@
DeepSeek-V3 $0.87_{0.80}^{0.93}$ $0.90_{0.83}^{0.93}$ $0.84_{0.80}^{0.90}$ $0.58_{0.53}^{0.63}$ $0.58_{0.53}^{0.63}$ $0.48_{0.43}^{0.57}$ $0.07_{0.03}^{0.10}$ $0.17_{0.07}^{0.23}$ $0.07_{0.00}^{0.17}$ $0.02_{0.00}^{0.03}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.48_{0.40}^{0.57}$ $0.71_{0.67}^{0.77}$ $0.59_{0.50}^{0.63}$ $0.62_{0.57}^{0.67}$ $0.61_{0.50}^{0.70}$ $0.22_{0.20}^{0.23}$ $0.29_{0.23}^{0.37}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.88_{0.83}^{0.90}$ $0.64_{0.57}^{0.73}$ $0.46_{0.37}^{0.53}$ $0.47_{0.33}^{0.53}$ $0.39_{0.30}^{0.43}$ $0.13_{0.10}^{0.17}$ $0.13_{0.10}^{0.17}$ $0.07_{0.00}^{0.13}$ $0.02_{0.00}^{0.03}$
o1-mini $0.62_{0.57}^{0.67}$ $0.70_{0.63}^{0.80}$ $0.27_{0.23}^{0.30}$ $0.18_{0.10}^{0.27}$ $0.14_{0.10}^{0.23}$ $0.10_{0.07}^{0.13}$ $0.03_{0.00}^{0.10}$ $0.06_{0.03}^{0.07}$ $0.02_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$
o3-mini $0.69_{0.60}^{0.80}$ $0.38_{0.23}^{0.47}$ $0.38_{0.33}^{0.47}$ $0.39_{0.23}^{0.50}$ $0.52_{0.47}^{0.60}$ $0.30_{0.23}^{0.37}$ $0.44_{0.33}^{0.50}$ $0.24_{0.20}^{0.27}$ $0.03_{0.00}^{0.07}$ $0.18_{0.13}^{0.23}$

### v2

@@ Appendix G Full Results over Problems
QwQ-32B $0.94_{0.90}^{1.00}$ $0.87_{0.83}^{0.93}$ $0.80_{0.70}^{0.93}$ $0.62_{0.57}^{0.67}$ $0.33_{0.23}^{0.40}$ $16_{0.10}^{0.20} 0.0$$3_{0.00}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.69_{0.67}^{0.73}$ $0.36_{0.27}^{0.40}$ $24_{0.13}^{0.40} 0.0$$9_{0.03}^{0.13} 0.0$$0_{0.00}^{0.00} 0.0$$1_{0.00}^{0.03} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o-mini $0.70_{0.67}^{0.73}$ $26_{0.13}^{0.40} 0.0$$9_{0.07}^{0.10} 0.0$$8_{0.03}^{0.13} 0.0$$1_{0.00}^{0.03} 0.0$$0_{0.00}^{0.00} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o $0.73_{0.73}^{0.73}$ $0.39_{0.33}^{0.43}$ $0.22_{0.17}^{0.27}$ $12_{0.07}^{0.20} 0.0$$9_{0.00}^{0.13} 0.0$$1_{0.00}^{0.03} 0.0$$6_{0.00}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.99_{0.97}^{1.00}$ $0.80_{0.70}^{0.90}$ $0.74_{0.67}^{0.83}$ $0.64_{0.53}^{0.77}$ $0.32_{0.17}^{0.50}$ $0.23_{0.20}^{0.27}$ $0.27_{0.20}^{0.33}$ $0.16_{0.10}^{0.27}$ $10_{0.10}^{0.10} 0.0$$02_{0.00}^{0.07}$
DeepSeek-V3 $0.83_{0.77}^{0.90}$ $0.44_{0.30}^{0.53}$ $0.14_{0.10}^{0.20}$ $16_{0.13}^{0.17} 0.0$$9_{0.00}^{0.17} 0.0$$6_{0.03}^{0.07} 0.0$$6_{0.00}^{0.10} 0.0$$6_{0.00}^{0.10} 0.0$$1_{0.00}^{0.03} 0.0$$01_{0.00}^{0.03}$
DeepSeek-V3-2503 $0.99_{0.97}^{1.00}$ $0.82_{0.77}^{0.90}$ $0.51_{0.50}^{0.53}$ $0.38_{0.27}^{0.53}$ $0.16_{0.07}^{0.27}$ $14_{0.13}^{0.17} 0.0$$09_{0.07}^{0.10}$ $10_{0.07}^{0.17} 0.0$$6_{0.03}^{0.07} 0.0$$03_{0.00}^{0.07}$
DeepSeek-R1 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.91_{0.83}^{0.97}$ $0.76_{0.63}^{0.93}$ $0.64_{0.57}^{0.73}$ $0.49_{0.43}^{0.57}$ $0.36_{0.27}^{0.40}$ $17_{0.13}^{0.23} 0.0$$04_{0.00}^{0.10}$
o1-mini $0.72_{0.57}^{0.80}$ $0.71_{0.67}^{0.77}$ $0.54_{0.50}^{0.60}$ $0.40_{0.33}^{0.47}$ $0.19_{0.17}^{0.23}$ $0.23_{0.13}^{0.30}$ $12_{0.10}^{0.17} 0.0$$8_{0.03}^{0.10} 0.0$$4_{0.00}^{0.07} 0.0$$00_{0.00}^{0.00}$
o3-mini $0.82_{0.80}^{0.83}$ $0.84_{0.77}^{0.90}$ $0.71_{0.63}^{0.77}$ $0.71_{0.60}^{0.83}$ $0.63_{0.57}^{0.73}$ $0.59_{0.50}^{0.67}$ $0.44_{0.37}^{0.50}$ $0.32_{0.27}^{0.43}$ $0_{0.10}^{0.33} 0.220$$2_{0.20}^{0.23}$
@@
QwQ-32B $0.88_{0.80}^{0.93}$ $0.83_{0.77}^{0.87}$ $46_{0.40}^{0.57} 0.0$$8_{0.00}^{0.20} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-R1-32B $26_{0.17}^{0.33} 0.0$$3_{0.00}^{0.07} 0.0$$1_{0.00}^{0.03} 0.0$$3_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o-mini $_{0.23}^{0.33} 0.030.03$$_{0.03}^{0.03}$ $04_{0.00}^{0.10} 0.0$$3_{0.00}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o $0.83_{0.73}^{0.90}$ $0.44_{0.40}^{0.50}$ $0.34_{0.33}^{0.37}$ $18_{0.13}^{0.20} 0.0$$4_{0.00}^{0.10} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.98_{0.97}^{1.00}$ $0.89_{0.83}^{0.93}$ $0.58_{0.43}^{0.70}$ $39_{0.37}^{0.40} 0.0$$7_{0.00}^{0.17} 0.0$$1_{0.00}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$3_{0.00}^{0.07} 0.0$$1_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3 $0.66_{0.60}^{0.73}$ $6_{0.40}^{0.50} 0.440$$4_{0.37}^{0.50}$ $37_{0.33}^{0.40} 0.0$$6_{0.00}^{0.13} 0.0$$4_{0.03}^{0.07} 0.0$$2_{0.00}^{0.03} 0.0$$0_{0.00}^{0.00} 0.0$$1_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.87_{0.80}^{0.93}$ $0.74_{0.67}^{0.83}$ $0.62_{0.57}^{0.67}$ $0.18_{0.10}^{0.27}$ $18_{0.13}^{0.23} 0.0$$9_{0.03}^{0.13} 0.0$$2_{0.00}^{0.07} 0.0$$2_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.80_{0.73}^{0.90}$ $0.64_{0.57}^{0.77}$ $0.49_{0.47}^{0.53}$ $29_{0.20}^{0.43} 0.0$$6_{0.03}^{0.10} 0.0$$03_{0.00}^{0.07}$
o1-mini $0.67_{0.43}^{0.80}$ $0.58_{0.50}^{0.63}$ $0.52_{0.47}^{0.57}$ $0.33_{0.27}^{0.40}$ $0.31_{0.20}^{0.50}$ $19_{0.13}^{0.23} 0.0$$7_{0.03}^{0.10} 0.0$$0_{0.00}^{0.00} 0.0$$2_{0.00}^{0.07} 0.0$$01_{0.00}^{0.03}$
@@
QwQ-32B $0.96_{0.90}^{1.00}$ $0.91_{0.87}^{0.93}$ $0.78_{0.73}^{0.87}$ $0.56_{0.50}^{0.67}$ $0.34_{0.27}^{0.43}$ $10_{0.07}^{0.13} 0.0$$1_{0.00}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-R1-32B $0.49_{0.43}^{0.57}$ $51_{0.43}^{0.57} 0.0$$3_{0.00}^{0.07} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o-mini $0.40_{0.30}^{0.50}$ $17_{0.13}^{0.20} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o $0.60_{0.57}^{0.63}$ $39_{0.30}^{0.53} 0.0$$3_{0.00}^{0.10} 0.0$$1_{0.00}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.76_{0.67}^{0.87}$ $0.70_{0.67}^{0.73}$ $0.22_{0.07}^{0.33}$ $17_{0.13}^{0.20} 0.0$$9_{0.07}^{0.10} 0.0$$4_{0.03}^{0.07} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3 $0.67_{0.63}^{0.70}$ $0.60_{0.53}^{0.63}$ $0.13_{0.07}^{0.20}$ $12_{0.10}^{0.17} 0.0$$3_{0.00}^{0.07} 0.0$$2_{0.00}^{0.07} 0.0$$2_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3-2503 $0.80_{0.73}^{0.87}$ $0.90_{0.87}^{0.93}$ $0.48_{0.40}^{0.63}$ $0.64_{0.57}^{0.73}$ $0.32_{0.30}^{0.37}$ $0.16_{0.07}^{0.20}$ $16_{0.07}^{0.23} 0.0$$9_{0.00}^{0.20} 0.0$$2_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1 $0.99_{0.97}^{1.00}$ $1.00_{1.00}^{1.00}$ $_{0.93}^{1.00} 0.970.97$$_{0.97}^{0.97}$ $0.88_{0.80}^{0.93}$ $0.72_{0.67}^{0.77}$ $0.72_{0.67}^{0.80}$ $0.51_{0.40}^{0.67}$ $22_{0.17}^{0.27} 0.0$$04_{0.03}^{0.07}$
o1-mini $0.61_{0.50}^{0.70}$ $0.76_{0.70}^{0.87}$ $0.57_{0.43}^{0.70}$ $0.62_{0.60}^{0.67}$ $0.37_{0.33}^{0.43}$ $0.27_{0.23}^{0.30}$ $0.34_{0.27}^{0.40}$ $17_{0.07}^{0.23} 0.0$$3_{0.00}^{0.07} 0.0$$02_{0.00}^{0.07}$
@@
DeepSeek-V3 $0.87_{0.80}^{0.93}$ $0.90_{0.83}^{0.93}$ $0.84_{0.80}^{0.90}$ $0.58_{0.53}^{0.63}$ $0.58_{0.53}^{0.63}$ $48_{0.43}^{0.57} 0.0$$07_{0.03}^{0.10}$ $17_{0.07}^{0.23} 0.0$$7_{0.00}^{0.17} 0.0$$02_{0.00}^{0.03}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.48_{0.40}^{0.57}$ $0.71_{0.67}^{0.77}$ $0.59_{0.50}^{0.63}$ $0.62_{0.57}^{0.67}$ $0.61_{0.50}^{0.70}$ $0.22_{0.20}^{0.23}$ $29_{0.23}^{0.37} 0.0$$2_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.88_{0.83}^{0.90}$ $0.64_{0.57}^{0.73}$ $0.46_{0.37}^{0.53}$ $0.47_{0.33}^{0.53}$ $0.39_{0.30}^{0.43}$ $0.13_{0.10}^{0.17}$ $13_{0.10}^{0.17} 0.0$$7_{0.00}^{0.13} 0.0$$02_{0.00}^{0.03}$
o1-mini $0.62_{0.57}^{0.67}$ $0.70_{0.63}^{0.80}$ $0.27_{0.23}^{0.30}$ $0.18_{0.10}^{0.27}$ $0.14_{0.10}^{0.23}$ $10_{0.07}^{0.13} 0.0$$3_{0.00}^{0.10} 0.0$$6_{0.03}^{0.07} 0.0$$2_{0.00}^{0.07} 0.0$$01_{0.00}^{0.03}$
o3-mini $0.69_{0.60}^{0.80}$ $0.38_{0.23}^{0.47}$ $0.38_{0.33}^{0.47}$ $0.39_{0.23}^{0.50}$ $0.52_{0.47}^{0.60}$ $0.30_{0.23}^{0.37}$ $0.44_{0.33}^{0.50}$ $24_{0.20}^{0.27} 0.0$$03_{0.00}^{0.07}$ $0.18_{0.13}^{0.23}$